three balls are drawn in succession without replacement from a jar containing 8 white balls and 5 black balls. Let B be the random variable representing number of black balls


1
Expert's answer
2022-05-09T16:21:42-0400

Since it is not stated which parameter has to be calculated, consider possible values that the random variable "B" may take. Since the number of balls are "3", the random variable "B" may take the following values: "0,1,2,3." The total number of balls is "8+5=13". There are "C_{13}^3=\frac{13!}{10!3!}=\frac{11\cdot12\cdot13}{3!}=286" different ways to choose "3" balls from "13". There are "C_8^{3}=\frac{8!}{5!3!}=\frac{6\cdot7\cdot8}{6}=56" different ways to choose "3" white balls from "8". We receive that: "P(B=0)=\frac{C_8^3}{C_{13}^3}=\frac{56}{286}\approx0.196". (it is rounded to "3" decimal places) There are "C_8^2=\frac{{8!}}{6!2!}=28" ways to choose "2" white balls from "8". There are "C_5^1=5" different ways to choose "1" black ball from "5". Using the multiplication principle of combinatorics, we get: "P(B=1)=\frac{C_8^2C_5^1}{C_{13}^3}=\frac{28\cdot 5}{286}\approx0490." We continue further without explanations and receive: "P(B=2)=\frac{C_8^1C_5^2}{C_{13}^3}=\frac{8\cdot10}{286}\approx0.280", "P(B=3)=\frac{C_5^3}{C_{13}^3}\approx0.035".

Answer: "P(B=0)\approx0.196", "P(B=1)\approx0.490", "P(B=2)\approx0.280", "P(B=3)\approx0.035".(all values are rounded to "3" decimal places)